# The Physics of Cometary Anti-tails as Observed in 3I/ATLAS

### v1

@@ Abstract
Observations of interstellar comet 3I/ATLAS at 3.8 au show an elongated coma similar to a cometary tail but pointing in the direction of the Sun. This type of anti-tail, not a result of perspective, may not have been previously observed. We explain the anti-tail as an anisotropic extension of the snow line, or survival radius of a sublimating ice grain, in the direction of the Sun. The anisotropy is due to the difference in the sublimation mass flux in the solar and perpendicular directions caused by the change in the illumination angle of the cometary surface. The stronger sublimation mass flux in the solar direction results in ice grains with larger sizes, longer sublimation lifetimes, and a snow line at a larger radial distance with respect to other directions. The observed radial surface brightness profiles as a function of illumination angle are well reproduced by a Haser-type spherical outflow with constant velocity and sublimating ice grains with angularly dependent survival lengths.
[…]
show]eketo@cfa.harvard.edu
show]aloeb@cfa.harvard.edu
@@ 1 Introduction
Observations of the interstellar object, 3I/ATLAS, by the Hubble Space Telescope (HST) on July 21, 2025 (D. Jewitt et al., 2025) revealed a faint anti-tail in the solar direction that is not due to projection. This phenomenon, observed at a distance of 3.8 au from the Sun, is not common and possibly observed for the first time in 3I/ATLAS. The anti-tail has been confirmed in other observations (J. Tonry et al., 2025; B. T. Bolin et al., 2025). In this paper, we explain the development of an anti-tail based on simple physical models of comets.
We model the coma as a spherical outflow of sublimating grains and reproduce the observed surface brightness profiles with the survival or snow-line length, the radial distance that an ice grain persists while sublimating, varying with illumination angle. Calculated surface temperatures and grain survival times indicate that the sublimation mass flux off the cometary surface must be predominantly CO2 as observed (C. M. Lisse et al., 2025; M. A. Cordiner et al., 2025), while the grains must be H2O ice.
From a chain of simple calculations, we determine the maximum ice grain size and terminal velocity as a function of the illumination angle to derive the survival length, $\ell=v_{\infty}t_{\rm life}$, the extent of the coma as a function of illumination angle.
@@ 2 The anti-tail in the HST image
Figure 1 shows two representations of the HST data. The image (left) is shown with a logarithmic color transfer function. Radial profiles (right) are shown along directions of 10, 100, 190 with respect to the coordinate axes. The profile of 10, in the direction of the Sun, is larger than the profiles in the two perpendicular directions at 100 and 190. The elongation of the coma toward the Sun may be called an anti-tail since it is in the opposite direction of a typical cometary tail. None of the profiles fits a single-slope power law. Their curvature suggests an exponential decrease in the number density of reflecting ice grains with distance from the nucleus as previously mentioned (D. Jewitt et al., 2025).
Refer to caption Refer to caption
@@
Figure 1: Left: HST image. Right: Radial profiles of the surface brightness at angles of 10, 100, and 190 where the angles are measured with respect to the $x,y$ coordinate axes. The observed profiles are overlaid with fits to a model Haser-type surface brightness. The $x,y$ axes have units of pixels (0.04”) and counts (electrons s-1).
@@ 3 A physical model for the anti-tail
We model the surface brightness profiles with a Haser-type density law (L. Haser et al., 2020) that develops from flux conservation in a constant-velocity, spherical outflow combined with destruction of the particles,
$\frac{1}{r^{2}}\frac{d}{dr}[r^{2}v(a)n(r,a)]=-\frac{n(r,a)}{t_{\rm life}}.$ (1)
where $a$ is the ice grain size with lifetime $t_{\rm life}$. The solution for the number density modifies the solution for a steady-state outflow by an exponential,
$n(r,a)\propto r^{-2}\exp\bigg{(}-\frac{r}{\ell(a)}\bigg{)}$ (2)
if we select a representative ice grain size to replace the ice-grain size distribution, we can use the same integral solution for the surface brightness as in the original solution of Haser developed for molecular emission. The surface brightness follows,
$\Sigma(\rho,a)\propto\frac{Q}{4\pi v(a)\rho}2K_{0}(x)$ (3)
where $\rho$ is the projected radius, $Q$ is the mass flux of the outflow, $v(a)$ is the constant terminal velocity, $K_{0}$ is the modified Bessel function of the second kind, and $x=\rho/\ell(a)$. The horizontal dashed line on figure 1 shows the approximate noise level. The vertical dashed line is placed at three times the resolution of the HST, about $3\times\ 2$ pixels. The fitting is restricted to data that lie above and to the right of the dashed lines.
The model curves that are fit to the data, give an empirical estimate for the lengths, $\ell$, of the snow lines in the different directions. At a heliocentric distance of 3.8 au, the apparent length in the solar direction is 3500 km, and 1300 km in the two perpendicular directions. At this distance, the comet is moving toward both the Sun and the Earth with a projection angle of 9.6 and a foreshortening factor of 5.8. Scaling the projected radius, $\rho$, implies a snow-line length in the solar direction of 29,600 km.
@@
The velocity of the ice grains is estimated with the Haser-Whipple (L. Haser, 1957; F. L. Whipple, 1951) model for free molecular acceleration,
@@
where $C_{D}$ is the drag coefficient ($\sim 1$); $v_{g}=\sqrt{8k_{B}T/(\pi m_{g})}$ ($\rm ms^{-1}$) is the gas velocity where $T<200$ K is the surface temperature ; and $J$ is the mass flux (kg $\rm m^{-2}s^{-1}$) from sublimation. Integration of the acceleration with some assumptions results in the Probstein terminal velocity (M. L. Finson & R. F. Probstein, 1968),
$v_{\infty}\approx\sqrt{\frac{3C_{D}}{2}}\sqrt{\frac{v_{g}R_{n}J}{\rho_{d}a}}\propto J^{1/2}{a^{-5/4}}$ (6)
where $R_{n}$ is the radius of the cometary nucleus. This velocity gives us the first of the two factors in the survival length scale $\ell(a)=v_{\infty}(a)t_{\rm life}$.
@@
Ice grain sizes are described by a power-law distribution with an index that is reliably maintained at 3.5-3.5 by fragmentation. The maximum ice grain size that can escape the gravitational pull of the nucleus is given by Whipple’s maximum liftable grain size. Although this maximum size grain will not actually escape the nucleus since its terminal velocity is zero, we can use this maximum as a multiplicative factor on the sizes of ice grains across the distribution. From the balance between gravity and the drag force.
@@
where $g$ is the gravitational force of the nucleus, $R_{n}$ is the nuclear radius, $\rho_{n}$ is nuclear density, and $J$ is the mass flux off the nucleus.
Since the terminal velocities scale as $v_{\infty}\propto a^{-5/4}$, and the ice grain lifetimes scale as $t_{\rm life}\propto a$, the survival length scales as $a^{-1/4}$. The ice grain sizes and hence their lifetimes and also their terminal velocities are both related to the mass flux off the nucleus, $J$.
The mass flux derives from the energy balance at the surface of the nucleus (D. Prialnik, 2004),
$(1-A)\frac{S_{0}}{r^{2}_{H}}\cos(\theta)=\epsilon\sigma T^{4}+L_{s}J-k\frac{\partial T}{\partial z}$ (9)
where $A$ is the Bond (bolometric) albedo; $S_{0}$ is the solar constant at 1 au (1361 W $\rm m^{-2}$); $r_{H}$ is the heliocentric distance in au; $\cos(\theta)$ is the illumination factor with $\theta$ the illumination or hour angle (HA) around the nuclear circumference and $\theta=0$ towards the Sun; $\epsilon$ is the thermal emissivity ($\sim 0.9$); and, $\L _{s}$ is the latent heat of sublimation.
The left side of the equation is the power from insolation. The first term on the right side is the radiative cooling. The second is the cooling from sublimation. The third term is the heat flow by conduction from the nuclear interior required to maintain a non-zero surface temperature on the dark side of the nucleus. The solution for the diffusion equation for heat flow is obtained with a standard conduction closure approximation that depends on the thermal inertia, $\Gamma=\sqrt{\kappa\rho c_{p}}\rm\ \ (J\ m^{-2}K^{-1}s^{-1/2})$ of the nucleus (M. Delbo et al., 2015).
@@
We also need the thermal inertia, $\Gamma$. If $\Gamma=25$ and $\Gamma=\sqrt{\kappa\rho c_{p}}$, $\rho_{n}=500\rm\ g\ m^{-3}$, and $c_{p}=500-1200\rm\ J\ kg^{-1}K^{-1}$, then $\kappa=(1.0-2.5)\times 10^{-}3\rm\ W\ m^{-1}K^{-1}$, reasonable numbers for comets (O. Groussin et al., 2019; J. Boissier et al., 2011).
Figure 2 shows the solutions for each of the pure species and for the mixture. The rapid sublimation of CO2 cools the surface of the mixture and keeps the temperature of the H2O too cool for sublimation as previously noted (C. M. Lisse et al., 2025). The gas coma is then dominated by CO2 as observed. Aside from the left-right symmetry, there is a 1-1 relationship between the temperature and the hour angle.
Refer to caption Refer to caption
Figure 2: (Left) Surface temperature of the nucleus as a function of angle for H2O, CO2, and mixed 80% H2O 20% CO2. (Right) Ratio $\psi$ (equation 16)
@@
The response time varies from a minimum of 170 s at HA=0 to 3.7 hr at HA=90.0 both shorter than the 16 hour observed rotation period (T. Santana-Ros et al., 2025; R. de la Fuente Marcos et al., 2025), validating the approximation of an equilibrium temperature.
@@
Figure 3 shows the length of the snow line as a function of hour angle for three different ice grain sizes. The maximum ice grain size, which has an outward velocity of zero by definition, represents the maximum ice grain size for the distribution of sizes and decreases with hour angle due to the diminishing sublimation mass flux. The factor $a_{\rm frac}$, shown in the legend represents a multiplier on the maximum ice grain size at each hour angle chosen to result in a representative ice grain size near the wavelength of optical light. Ice grains of this size dominate the total scattering cross-section. The snow line, $\ell(a,T_{s})$ is longest toward the Sun and falls off with hour angle and surface temperature. Therefore, the coma should be extended toward the Sun as observed.
The curves for the snow line have the same functional dependence on hour angle as the maximum ice grain size and the sublimation mass flux of CO2. All are well-fit by a cosine function implicating the illumination angle in the energy balance as the dominant effect on the length of the snow line.

### v2

@@ Abstract
Observations of interstellar comet 3I/ATLAS at 3.8 au show an elongated coma similar to a cometary tail but pointing in the direction of the Sun, possibly the first instance of this type of anti-tail which is not a result of perspective. We explain the anti-tail as an anisotropic extension of the snow line, or survival radius of a sublimating ice grain, in the direction of the Sun. The anisotropy is due to the difference in the sublimation mass flux in the solar and perpendicular directions caused by the change in the illumination angle of the cometary surface. The stronger sublimation mass flux in the solar direction results in ice grains with larger sizes, longer sublimation lifetimes, and a snow line at a larger radial distance with respect to other directions. The observed radial surface brightness profiles as a function of illumination angle are well reproduced by a Haser-type radial outflow with constant velocity and sublimating ice grains with angularly dependent survival lengths.
keywords:
comets: general, comets: individual: 3I/ATLAS
pubyear: 2025pagerange: The Physics of Cometary Anti-tails as Observed in 3I/ATLASData Availability
@@ 1 Introduction
Observations of the interstellar object, 3I/ATLAS, at a distance of 3.8 au from the Sun, by the Hubble Space Telescope (HST) on July 21, 2025 (Jewitt_2025) revealed an anti-tail in the solar direction that is not due to projection. Confirmed by other observations (Tonry_2025; Bolin_2025; Cordiner_2025; Chandler_2025; Seligman_2025), this phenomenon is not common and possibly observed for the first time. In this paper, we explain the development of an anti-tail based on simple physical models of comets.
We model the coma as a radial outflow of sublimating ice grains ejected by gas drag off the cometary surface. Calculated surface temperatures and grain survival times indicate that the sublimation mass flux off the cometary surface must be predominantly CO2 as observed (Lisse_2025; Cordiner_2025), while the grains must be H2O ice. The radial extent of the coma depends on the distance that an ice grain survives while sublimating. This snow-line length varies with the solar illumination angle principally through the sublimation mass flux which determines the maximum liftable grain size (Whipple_1951) and the maximum of the grain size distribution. From a chain of simple calculations, we determine the maximum ice-grain size and terminal velocity as a function of the illumination angle to derive the survival length, $\ell=v_{\infty}t_{\rm life}$ as a function of illumination angle.
@@ 2 The anti-tail in the HST image
Figures 1 and 2 show two representations of the HST data. The image is shown with a logarithmic color transfer function. Radial profiles are shown along directions of 10, 100, 190 with respect to the coordinate axes. The profile of 10, in the direction of the Sun, is longer than the profiles in the two perpendicular directions at 100 and 190. None of the profiles fits a single-slope power law as would be expected from a radial dilution of the number density in a constant-velocity, spherical outflow. Their curvature suggests an exponential decrease in the number density with distance from the nucleus as previously mentioned (Jewitt_2025).
We can model the curvature of the radial surface brightness profiles with a Haser-type density law (Haser_2020) that develops from flux conservation in a constant-velocity, spherical outflow combined with destruction of the particles,
$\frac{1}{r^{2}}\frac{d}{dr}[r^{2}v(a)n(r,a)]=-\frac{n(r,a)}{t_{\rm life}}.$ (1)
where $a$ is the ice grain size with velocity, $v(a)$ and lifetime $t_{\rm life}$. The solution for the number density modifies the solution for a steady-state outflow by an exponential,
$n(r,a)\propto r^{-2}\exp\bigg(-\frac{r}{\ell(a)}\bigg).$ (2)
If we select a representative ice grain size to replace the ice-grain size distribution, we can use the same integral solution for the surface brightness as in the original Haser solution for molecular emission. The surface brightness follows,
$\Sigma(\rho,a)\propto\frac{Q}{4\pi v(a)\rho}2K_{0}(x)$ (3)
where $\rho$ is the projected radius, $Q$ is the mass flux of the outflow and $K_{0}$ is the modified Bessel function of the second kind, and $x=\rho/\ell(a)$.
Figure 2 shows the mathematical fits to the Haser model,
$I(x)=C+(A/x)\exp{(x/B)}$ (4)
where $I(x)$ is the observed counts at coordinate $x$ (pixels) and $A,B,C$ are constants. The fitting is restricted to data that lie above and to the right of the dashed lines. The horizontal dashed line on figure 1 shows the approximate noise level. The vertical dashed line is placed at three times the resolution of the HST, about $3\times\ 2$ pixels.
The model curves give an empirical estimate for the lengths, $\ell$, of the snow lines in the different directions. At a heliocentric distance of 3.8 au, the apparent length in the solar direction is 3500 km, and 1300 km in the two perpendicular directions. At this distance, the comet is moving toward both the Sun and the Earth with a projection angle of 9.6 and a foreshortening factor of 5.8. Scaling the projected radius, $\rho$, implies a snow-line length in the solar direction of 29,600 km. The characteristic length of the exponential destruction indicates that volatile ice grains, rather than refractory dust grains, contribute most of the total scattering cross-section in the coma.
Figure 1: Left: HST image in log10 counts (electrons s-1).
@@
Figure 2: Radial profiles of the surface brightness at angles of 10, 100, and 190 where the angles are measured with respect to the $x,y$ coordinate axes. The observed profiles are overlaid with fits to a model Haser-type surface brightness. The $x,y$ axes have units of pixels (0.04") and counts (electrons s-1).
The HST observations indicate that the sunward anti-tail from 3I/ATLAS is different from the two more commonly observed anti-tails. Ejection of gas and dust is commonly anisotropic in the direction of the Sun, and particularly active regions of the nucleus can produce jet-like structures directed toward the Sun. Second, the projection of the current position of a coma against a stream of previously ejected dust may place the stream on the sunward side of the coma due to the difference in viewing perspective at the past time of ejection and the current time.
Sunward-directed jets are characteristically narrower, shorter, and fainter than the comae in which they are embedded. Although they can create asymmetry in the larger comae, they may not be easily apparent without image enhancement, for example with the Larson-Sekanina rotational-gradient method of difference imaging (Larson_1984). This filtering is scale dependent according to the spatial separations, ($\Delta r,\Delta\alpha)$, in their equation 1,
$\begin{split}B^{\prime}(r,\alpha;\Delta r,\Delta\alpha)&=[B(r,\alpha)-B(r-\Delta r,\alpha-\Delta\alpha)]\\ &\quad+[B(x,P)-B(x-\Delta x,P+\Delta P)]\end{split}$ (5)
where $B$ is the image brightness in polar coordinates ($r,\alpha)$. Application to the HST image, shown in figure 3, finds no anisoptropies on the scale of $(\Delta r,\Delta\alpha)=(3,10)$ (pixels, degrees) as might be expected for a jet-like ejection. The broader anisoptropy of the 3I/ATLAS anti-tail begins to show in the differenced image on scales of $(\Delta x,\Delta)\alpha\sim(60,30)$. This is consistent with the hypothesis that the 3I/ATLAS anti-tail is created by anisotropic destruction by sublimation of the ice grains rather than anisotropic jet emission.
Figure 3: HST image filtered by the rotational-gradient transform with ($\Delta r,\Delta\alpha)=(3,10)$ (equation 5 showing an absence of small scale anisotropy.
Apparent anti-tails due to projection are sometimes seen when a periodic comet crosses its aphelion orbital point (opposition or true anomaly of 180). At this point the comet’s orbital plane may be seen nearly edge-on. Large dust grains ejected at the comet’s previous perihelion passage or passages, whose orbits approximate the comet’s, may appear in projection with the comet as narrow sunward tails or necklines. In contrast, 3I/ATLAS is on the incoming leg of its hyperbolic orbit. Furthermore, the exponential decline in the number density of scatterers indicated by the Haser-type surface brightness profiles suggests that the scatterers are destroyed before they have time to wrap around the comet to form a traditional antisolar tail whose remnants might be seen in projection with the coma.
For further comparison, the images of the comet 67P / Churyumov-Gerasimenko analyzed in Rosenbush_2026 illustrate the three types of previously known tails, a traditional dust tail, jet-like ejection revealed by the rotational-gradient filter, and projected necklines. In the 6 October 2021 observations (Fig. 3), comet 67P shows a normal antisolar dust tail (PA $\approx 266^{\circ}$) and a sunward jet (PA $\approx 127^{\circ}$) produced by an active area on the nucleus. In the later 6 February 2022 images (Fig. 4), a third feature appears, a bright sunward linear structure identified as a neckline, composed of old dust ejected near perihelion that follows the comet but on a slightly different orbital path. At the same time, two active jets (J1, J2) are seen roughly perpendicular to the Sun-comet vector, representing current outgassing activity. The 3I/ATLAS anti-tail does not fall easily into any one of these categories.
This discussion in this section indicates that the anti-tail from 3I/ATLAS develops from the dependence of the length of the snow line on the illumination angle. So far, we have a mathematical fit for the radial surface brightness in terms of the Haser model that provides estimates of the lengths but not a description of the physics. We now proceed to estimate the snow-line length from physical principles which is the main goal of this study.
@@ 3 A physical model for the anti-tail
The velocity of the ice grains is estimated with the Haser-Whipple (Haser_1957; Whipple_1951) model for free molecular acceleration,
@@
where $C_{D}\sim 1$ is the drag coefficient; $v_{g}=\sqrt{8k_{B}T/(\pi m_{g})}$ ($\rm m\ s^{-1}$) is the free-expansion gas velocity where $T<200$ K is the surface temperature ; and $J$ is the mass flux (kg $\rm m^{-2}s^{-1}$) from sublimation; and $\rho_{d}$ is the grain density. Integration of the acceleration with some assumptions results in the Probstein terminal velocity (Finson_1968),
$v_{\infty}\approx\mathrm{min}\bigg[\sqrt{\frac{3C_{D}}{2}}\sqrt{\frac{v_{g}R_{n}J}{\rho_{d}a}},v_{\mathrm{g}}\bigg].$ (8)
The radius of the cometary nucleus, $R_{n}$, is used as a characteristic length for the acceleration. This estimate requires capping the terminal velocity at the free-expansion speed of the gas, $v_{\mathrm{g}}$. The terminal velocity scales as $v_{\infty}\propto J^{1/2}a^{-5/4}$ up to its maximum. This velocity gives us the first of the two factors in the survival length scale $\ell(a)=v_{\infty}(a)t_{\rm life}$.
@@
Ice grain sizes are described by a power-law distribution with an index that is reliably maintained at 3.5-3.5 by fragmentation. The maximum ice grain size that can escape the gravitational pull of the nucleus is given by Whipple’s maximum liftable grain size. Although this maximum size grain will not actually escape the nucleus since its terminal velocity is zero, we can use this maximum of the grain-size distribution. From the balance between gravity and the drag force.
@@
where $g$ is the gravitational force of the nucleus, $R_{n}$ is the nuclear radius, $\rho_{n}$ is nuclear density, and $J$ is the mass flux off the nucleus. The survival length of larger grains whose terminal velocities are below the maximum of the free-expansion velocity scales as $a^{-1/4}$. For smaller grains that are accelerated to $v_{\mathrm{g}}$, the maximum velocity, $\ell\propto a$.
The ice grain sizes and hence their lifetimes and also their terminal velocities are both related to the mass flux off the nucleus, $J$.
The mass flux derives from the energy balance at the surface of the nucleus (Prialnik_2004),
$(1-A)\frac{S_{0}}{r^{2}_{H}}\cos(\theta)=\epsilon\sigma T^{4}+L_{s}J-\Gamma\sqrt{\omega/\pi}~(T-T_{b})$ (11)
where $A$ is the Bond (bolometric) albedo; $S_{0}$ is the solar constant at 1 au (1361 W $\rm m^{-2}$); $r_{H}$ is the heliocentric distance in au; $\cos(\theta)$ is the illumination factor with $\theta$ the illumination or hour angle (HA) around the nuclear circumference and $\theta=0$ towards the Sun; $L_{s}$ is the latent heat of sublimation; $\omega$ is the angular frequency $2\pi/P$ for period $P$; and $T_{b}$ is the mean equilibrium temperature. $T_{b}$ is found from the same equation by dropping the last term, replacing $\cos\theta$ by its mean $\langle\cos\theta\rangle\approx 1/\pi$, and solving for $T$.
The left side of the equation is the power from insolation. The first term on the right side is the radiative cooling. The second is the cooling by sublimation. The third term is the heat flow by conduction from the nuclear interior. This maintains a minimum temperature on the dark side of the nucleus. The form in equation 11 is an approximation of $\kappa{\partial T}/{\partial z}$ for radial direction, $z$. The thermal inertia, $\Gamma$, and the thermal conductivity, $\kappa$ are related as $\Gamma=\sqrt{\kappa\rho_{n}c_{p}}\rm\ \ =10-60\ \ (J\ m^{-2}\ K^{-1}\ s^{-1/2})$, where $\rho_{n}$ is the nuclear density (Delbo_2015; Groussin_2019; Groussin_2025; Boissier_2011).
@@
Figure 4 shows the solutions for each of the pure species and for the mixture. The rapid sublimation of CO2 cools the surface of the mixture and keeps the temperature of the H2O too cold for sublimation as previously noted (Lisse_2025). The gas in the coma is then predominantly CO2 as observed (Lisse_2025; Cordiner_2025). Since the surface temperatures are too low for a significant sublimation mass flux of H2O, the relatively low proportion of gas phase H2O relatve to CO2 in the coma (Cordiner_2025) may derive from the sublimation of the ice grains ejected by gas drag.
Figure 4: Surface temperature of the nucleus as a function of angle for H2O, CO2, and mixed 80% H2O 20% CO2.
Figure 5: Ratio $\psi$ (equation 19)
@@
The response time varies from a minimum of 170 s at HA=0 to 3.7 hr at HA=90.0 both shorter than the 16 hour observed rotation period (SantanaRos_2025; FuenteMarcos_2025).
The thermal response time scale is longer at lower temperatures. Because of the thermal lag, the temperature on the dark side of the comet will be warmer than its equilibrium temperature We can estimate the surface temperatures more precisely by including the thermal lag. Written as a finite difference approximation, we estimate the temperature of an equatorial patch of the surface as it rotates with the comet, assuming for convenience a rotation axis perpendicular to the Sun-comet vector (maximum rotation speed),
$T_{n+1}=T_{b}(HA_{n+1})+[T_{n}-T_{\mathrm{eq}}(HA_{n+1})]e^{\Delta t/\tau_{n+1}}$ (15)
The time series converges immediately to a steady state and is shown in figure 6 for $\Gamma=25$. For comparison, figure 7 shows the result for $\Gamma=100$. The difference amounts to an increase in the minimum temperature of  7% on the dark side of the comet.
Figure 6: Equilibrium and time-dependent surface temperatures of an equatorial patch of the mixed composition nucleus during two rotations. The thermal inertia, $\Gamma=25$ J m-2 K-1 s-1/2.
Figure 7: Same as figure 6 but with the thermal inertia, $\Gamma=100$ J m-2 K-1 s-1/2.
@@
Figure 8 shows the length of a representative snow line as a function of hour angle. The figure shows that a representative snow line varies in length by a factor of about seven over the range of hour angle $\pm 90^{\circ}$. This variation creates the sunward anti-tail.
[…]
Aside from the asymmetry produced by the thermal lag, the curves for the snow line have the same functional dependence on hour angle as the maximum ice grain size and the sublimation mass flux of CO2. All are well-fit by a cosine function implicating the illumination angle in the energy balance as the dominant effect on the length of the snow line.
@@
4 Discussion
The model of the anti-tail as a result of the snow-line length in a coma composed of CO2 gas, H2O ice grains is an alternative to other hypotheses such as an anti-tail from enhanced ejection of large dust grains followed by fragmentation (Cordiner_2025); from sunward jets (Jewitt_2025b) and/or shaped by solar radiation pressure (Jewitt_2025); or, from anisotropic dust emission from a nucleus with an in-plane rotation pole (Chandler_2025). The snow-line model aligns with other observations. In particular, the surface temperature of the nucleus maintained by the sublimation of CO2 prevents significant sublimation of H2O off the nucleus. Instead, the H2O gas in the coma derives from the sublimation of the ice grains. This allows a comet with a volatile composition of 4:1 H2O:CO2 to produce a coma with a 1:8 composition in the gas phase (Keto_2025b) as observed by the JWST (Cordiner_2025). The snow-line model also allows for the disappearance of the anti-tail at heliocentric distances within $\sim 3.5$ au where the ice coma collapses due to higher insolation and rapid destruction of the ice grains (Keto_2025b) as observed (Hoogendam_2025; NOIRLab2525) .